Eliminate This Problem
x + y = 10
x - y = - 6

Answer by jim_thompson5910(35256)

x + y = 10
x - y = - 6
---------------
2x+0y = 4

2x = 4

x = 4/2

x = 2
=======================================================

x+y = 10

2+y = 10

y = 10-2

y = 8

So the answers are x = 2 and y = 8